The body with mass m = 1.2kg is pulled with the horizontal force F = 20 N on a rough horizontal surface and during 3 s makes the way 6 m starting from the resting position. Calculate the force and coefficient of sliding friction.
1
Expert's answer
2020-12-14T07:29:00-0500

The net force, acting on the body is "\bold F + \bold N + \bold F_f + m\bold g", and is equal to "m \bold a", by second Newton's law. In projections:

"oy": "N = mg",

"ox": "ma = F - F_f".

Kinetic friction is "F_f = \mu N", and using "oy" equation, "F_f = \mu m g". Substituting into "ox" equation, obtain "ma = F - \mu mg", or "a = \frac{F}{m} - \mu g".

Displacement of the block in time "t" is "s = \frac{a t^2}{2}", from where "a = \frac{2 s}{t^2} = \frac{F}{m} - \mu g".

From the last equation, coefficient of friction is "\mu = \frac{1}{g}\left[\frac{F}{m} - \frac{2 s}{t^2}\right] \approx 1.56", and "F_f = \mu m g = 18.4 N".